3c. The retention times tr for compounds A and B are 7 min and 5.5 min, respectively. Knowing that the column void volume is V0 = 1 ml and that the solvent flow rate for compound A is rA = 1 ml/min and for compound B is rB = 2 ml/min, calculate the retention (or capacity) factors for compounds A and B.

3d. Calculate the selectivity factor α for compounds A and B above and indicate if the two compounds can be well separated in the experimental conditions specified in question 3.c above.
Answer #1

3c. The retention factor for compound A (kA) = (tr) of A * rA/V0 = 7 min * (1 mL/min)/1 mL = 7

The retention factor for compound B (kB) = (tr) of B * rB/V0 = 5.5 min * (2 mL/min)/1 mL = 11

3d. The selectivity factor for the compounds A and B (α) = kB/kA = 11/7 = 1.57

Therefore, the the two compounds can be well separated in the experimental conditions specified in question 3c.